Integrate tan³x dx
1
Expert's answer
2020-02-17T12:05:23-0500

"\int \tan^3x\ dx ="


"\int \cfrac{\sin^2x\ \sin x}{\cos^3x}\ dx ="


"\int \cfrac{1 - \cos^2x}{\cos^3x} \sin x\ dx ="


"\int \cfrac{\cos^2x - 1}{\cos^3x}\ d(\cos x) ="


"\int \cfrac{d(\cos x)}{\cos x} - \int \cfrac{d(\cos x)}{\cos^3x} ="


"\ln|\cos x| + \cfrac1{2\cos^2x} + C"